# Video Understanding with Large Language Models: A Survey

### v7

@@ III-B1 Training-free Vid-LLMs
InternVideo2 [224] 8 InternVL-6B+VideoMAE-g Q-former + LoRA 256 A100 GPUs Mixer-7B + BERT-large 7B 03/2024
LangRepo [160] - - - - Mistral 7B/12B 03/2024
@@
LITA [253] 100 CLIP ViT-L/14 MLP + SlowFast Token Pooling 8 A100 GPUs Vicuna 7B/13B 03/2024
MovieLLM [225] 1 FPS CLIP ViT-L/14 - 4 A100 GPUs Vicuna 7B 03/2024
@@
OmniViD [254] 32 VideoSwin Mixed Q-former+Visual Translator 8 A100 GPUs GPT-4 - 03/2024
RAVA [176] 30 FPS - - - GPT-4 - 03/2024
SCHEMA [177] - CLIP ViT-L/14 - 1 V100 GPU GPT-3.5 - 03/2024
TV-TREES [179] 2 FPS CLIP ViT-L/14 - - GPT-3.5 - 03/2024
IG-VLM [159] 6 CLIP ViT-L/14 - - LLaVA-v1.6, and GPT-4V 7B/13B/34B 03/2024
AVicuna [120] 100 CLIP ViT MLP, LoRA 1 A6000 GPU Vicuna-v1.5 7B 03/2024
CAT [222] - ImageBind Linear Projector 1 A100 GPU LLaMA-2 7B 03/2024
Uni-AD [272] - CLIP ViT-B/32 Video mapping network 8 A100 GPUs LLaMA-2 7B 03/2024
Elysium [251] Varying CLIP-ViT-L T-Selector 24 A100-80G Vicuna - 03/2024
@@
VideoTree [173] Varying EVA-CLIP ViT-G/14 - 4 A6000 GPUs GPT-4 - 04/2024
VTG-LLM [259] 96 EVA-CLIP ViT-G/14 Projector 6 ATN 910B LLaMA-2 7B 04/2024
AutoAD III [208] 8 EVA-CLIP Linear Projector 1 A40 GPU GPT-3.5-turbo - 04/2024
LLaVA-Hound-DPO [210] 10 LanguageBind ViT/14 - - GPT-4V - 04/2024
RedViLLM [211] Varying CLIP ViT-G/14 Temporal Module - Qwen-VL 7B 04/2024
LAVAD [174] 16 BLIP-2 ViT-L/14 - - LLaMA-2-chat 13B 04/2024
@@
VLM4HOI [250] - EVA CLIP Projection Layer 8 V100 GPUs LLaMA-2-Chat 13B 04/2024
@@
Koala [212] 64 EVA-CLIP ViT-G/14 Video QFormer+Linear Layer 4 RTX A6000 GPUs Vicuna-v0 7B/13B 04/2024
LongVLM [213] 100 CLIP ViT-L/14 Projection Layer 4 A100 GPUs Vicuna-v1.1 7B 04/2024
MA-LMM [214] 100 EVA-CLIP ViT-G/14 Q-Former 4 A100 GPUs Vicuna 7B 04/2024
@@
MiniGPT4-Video [215] Varying EVA-CLIP Projector + LoRA for LLM - LlaMA-2/Mistral 7B 04/2024
MoReVQA [158] 1 FPS - - - PaLM-2 - 04/2024
Pegasus-v1 [216] - Marengo 2.6 VL Alignment Module - GPT-4 - 04/2024
PLLaVA [217] 16 CLIP ViT-L/14 MM Projector - LLaVA 7B/13B/34B 04/2024
ST-LLM [218] Varying CLIP ViT-L/14 Projection Layer 8 A100 GPUs Vicuna-v1.1 7B 04/2024
Tarsier [220] Varying CLIP ViT MLP 48 A100 GPUs Vicuna 7B/13B 04/2024
TraveLER [175] Varying - - 8 RTX A6000 GPUs GPT-3.5 - 04/2024
V2Xum-LLaMA [100] 1 FPS CLIP ViT-L/14 Vision Adapter 8 A100 GPUs LLaMA 7B/13B 04/2024
VITRON [260] - - Projection Layer 10 A100 GPUs Vicuna-v1.5 7B 04/2024
X-VARS [221] 16 CLIP ViT-L/14 Linear Layer 1 V100 GPU VideoChatGPT - 04/2024
VideoChat2 [203] Varying UMT-L QFormer - Vicuna 7B 05/2024
@@
Shotluck Holmes [204] 120 SigLip MLP Projection Layer 8 H100 GPUs Vicuna 7B 05/2024
VideoStreaming [205] 16 CLIP ViT-L/14 Projecter 32 A100 GPUs Vicuna 7B 05/2024
VideoNarrator [206] - CLIP ViT-L/14 Video Projecter 4 A6000 GPUs Baichuan 7B 05/2024
TOPA [207] 10 CLIP ViT-L/14 Linear Projector 4 A100 GPUs Llama2 7/8/13B 05/2024
MotionLLM [202] 8 LanguageBind/VQ-VAE Linear Layer A100 GPUs Vicuna 7B 05/2024
Artemis [191] Varying CLIP ViT-L/14 MLP 8 A800 GPUs Vicuna 7B 06/2024
DrVideo [169] 2 FPS CLIP ViT-L/14 - - LaViLa 7B 06/2024
@@
EmoLLM [192] 8 CLIP ViT-L/14 Multi-Perspective Visual Projection 4 4090 GPUs Vicuna-v1.5 7B 06/2024
@@
FTFV-LLM [193] 8 OpenCLIP ViT-G/14 Vision-Language Adapter 64 A100 GPUs Vicuna 7B 06/2024
Flash-VStream [194] 8 - MLP 8 A100 GPUs GPT-3.5 - 06/2024
@@
Holmes-VAD [248] Varying CLIP ViT-L/14 Temporal Sampler 2 A100 GPUs Vicuna-v1.5 7B 06/2024
LongVA [196] 1 FPS CLIP ViT-L/14 MLP 8 A100 GPUs Mistral-7B-Instruct-v0.2 7B 06/2024
OmAgent [170] Varying - - - GPT-4o - 06/2024
GIT-LLaVA [156] 6 CLIP ViT-L/14 MLP 4 A100 GPUs Vicuna 7B 06/2024
@@
ShareGPT4Video [197] 16 CLIP ViT-L/14 Linear Layer + LoRA for LLM 8 A100 GPUs LLaMA-3 8B 06/2024
LVNet [171] - - - - GPT-4 - 06/2024
VIM [198] - EVA-CLIP ViT-G/14 Q-former+LoRA in LLM 8 A100 GPUs Vicuna-v1.1 13B 06/2024
Video-SALMONN [199] 2 FPS BLIP-2 ViT-G/14 MRC Q-Former - GPT-4 - 06/2024
VideoGPT+ [200] 16/8 CLIP ViT-L/14 Linear Layer 8 A100 GPUs Phi-3-Mini 3.8B 06/2024
Vriptor [268] - EVA-CLIP ViT-G/14 - A100 GPUs - - 06/2024
MM-Screenplayer [157] - - - - GPT-4-turbo - 06/2024
VideoLLaMA 2 [201] 16 CLIP ViT-L/14 Linear Layer +STC connector - LLaVA-1.5 7B 06/2024
@@
VideoLLM-online [249] 2 FPS CLIP ViT-L/14 MLP projector + LoRA for LLM - Llama-2-Chat/Llama-3-Instruct 7B/8B 06/2024
@@ III-B2 Fine-tuning Vid-LLMs
  • LLM Fully Fine-tuning: This fine-tuning method does not use any adapters but instead employs supervised training with a lower learning rate, updating all the parameters in the LLM. This method allows the Vid-LLM to fully adapt to the respective task and achieve good performance, especially when the target task is quite different from the pretraining tasks. For end-to-end Vid-LLMs, especially those in the Video Embedder $\times$ LLM category, the Video Embedder may also be fine-tuned for more comprehensive learning. However, this method consumes more computational resources than adapter-based fine-tuning methods and may potentially impair the inherent capabilities of the LLM, such as zero-shot and in-context learning. Vid-LLM adopted LLM Fully Fine-tuning include AV-LLM [231] and Vid2Seq [271]. In [231], there are both fully fine-tuning and adapter fine-tuning versions of Vid-LLMs, and the former’s performance is better than the latter.
  • Connective Adapter Fine-tuning: Here, the term “Connective” refers to adapters that bridge the Video Embedder and the LLM externally, enabling information from the video to flow into the LLM through the Connective Adapter. As illustrated in Figure 6, during training, the parameters of both the Video Embedder and the LLM are frozen, and only the parameters of the Connective Adapter are updated. Common Connective Adapters include MLP/Linear Layer and Q-former [283], their combinations, etc., whose primary function is to map video embeddings from the visual semantic space to the text semantic space of the LLM input tokens (i.e., modality alignment). Typically, fine-tuning only the Connective Adapter does not alter the LLM’s inherent behavior.
@@ IV-A Closed-ended Evaluation
Closed-ended evaluations use pre-defined answers or structured formats [284]. These include multiple-choice questions (TVQA [109], How2QA [285], STAR [286]) and questions with structured formats for direct comparison with ground truth (MSRVTT-QA [93], MSVD-QA [93], MVBench [203]). Multiple-choice performance is evaluated via accuracy percentages, while structured formats use metrics like CIDEr [287], METEOR [288], ROUGE [289], and SPICE [290] to compare predictions with ground truth. Notable benchmarks include MSRVTT-QA [93], TVQA [109], MVBench [203], EgoSchema [291], and Video-MME [292]. Each targets different video understanding aspects: TGIF-QA [107] focuses on action recognition, state transition, frame-level QA, and counting; ActivityNet-QA [110] covers motion, spatial, temporal, and free-form dimensions; VidComposition [293] emphasizes compositional reasoning; while NExT-QA [111] and MLVU [294] include causal and temporal action reasoning. These diverse question types test various reasoning abilities, though many benchmarks still exhibit domain biases toward common scenarios and lack diversity in rare events or unusual contexts.
@@
MSRVTT-QA [93] 2,990 2,990 15.2 Open-domain 72,821 Closed-ended & open-ended (what/who/how/when/where)
MSVD-QA [93] 504 504 9.8 Open-domain 50,505 Closed-ended & open-ended (what/who/how/when/where)
TGIF-QA [107] 9,575 9,575 3.0 Open-domain 8,506 Closed-ended & open-ended, action, transition, counting
ActivityNet-QA [110] 800 800 111.4 Human activity 8,000 Closed-ended & open-ended (what/who/how/when/where/why)
TVQA [109] 2,179 15,253 11.2 TV show 15,253 Closed-ended, multiple-choice
How2QA [285] 1,166 2,852 15.3 TV episode 2,852 Multiple-choice
STAR [286] 914 7,098 11.9 Human action 7,098 Multiple-choice
NExT-QA [111] 1,000 1,000 39.5 Daily life 8,564 Multiple-choice and open-ended (causal, temporal, descriptive)
MVBench [203] 3,641 3,641 16.0 Open-domain 4,000 Closed-ended (various tasks)
Video-Bench [295] 5,917 5,917 56.0 Open-domain 17,036 Open-ended (various tasks)
EgoSchema [291] 5,063 5,063 180.0 Egocentric activity 5,063 Closed-ended, procedural understanding
AutoEval-Video [296] 327 327 14.6 Open-domain 327 Open-ended evaluation
TempCompass [297] 410 500 11.4 Open-domain 7,540 Both closed-ended & open-ended, temporal reasoning
Video-MME [292] 900 900 1,017.9 Open-domain 2,700 Closed-ended evaluation
VideoVista [298] 894 3,400 131.0 Open-domain 25,000 Open-ended (descriptive, causal, predictive)
CinePile [299] 9,396 9,396 160.0 Movie 303,828 Movie understanding, open-ended
SOK-Bench [300] 10,000 10,000 - Open-domain 44,000 Open-ended, subject-oriented knowledge
SFD [301] 1,078 1,078 780.0 Movies 4,885 Multiple-choice, open-ended
EditVid-QA [302] 32,000 32,000 - Entertainment 204,000 Open-ended, editing techniques
InfiniBench [303] 1,219 1,219 4,460.4 Movie/TV show 108,200 Closed-ended & open-ended, long-form understanding
MLVU [294] 1,334 1,334 180.0-7,200.0 Open-domain 2,593 Both closed-ended and open-ended, long-form understanding
MMWorld [304] 1,910 1,910 102.3 Open-domain 1,599 Open-ended, multimodal evaluation
VELOCITI [305] 900 900 10.0 Movie - Open-ended, visual understanding
VidComposition [293] 982 982 - Movie/TV show 1,706 Multiple-choice, compositional reasoning
EAGLE [306] 36,000 36,000 16.0-76.0 Egocentric activity 400,000 Open-ended, procedural understanding

### v8

@@ III-B1 Training-free Vid-LLMs
InternVideo2 [Wang2024InternVideo2] 8 InternVL-6B+VideoMAE-g Q-former + LoRA 256 A100 GPUs Mixer-7B + BERT-large 7B 03/2024
LangRepo [Kahatapitiya2024LangRepo] - - - - Mistral 7B/12B 03/2024
@@
LITA [Huang2024LITA] 100 CLIP ViT-L/14 MLP + SlowFast Token Pooling 8 A100 GPUs Vicuna 7B/13B 03/2024
MovieLLM [Song2024MovieLLM] 1 FPS CLIP ViT-L/14 - 4 A100 GPUs Vicuna 7B 03/2024
@@
OmniViD [Wang2024OmniViD] 32 VideoSwin Mixed Q-former+Visual Translator 8 A100 GPUs GPT-4 - 03/2024
RAVA [Cao2024RAVA] 30 FPS - - - GPT-4 - 03/2024
SCHEMA [Niu2024SCHEMA] - CLIP ViT-L/14 - 1 V100 GPU GPT-3.5 - 03/2024
TV-TREES [S2024TV-TREES] 2 FPS CLIP ViT-L/14 - - GPT-3.5 - 03/2024
IG-VLM [Kim2024IG-VLM] 6 CLIP ViT-L/14 - - LLaVA-v1.6, and GPT-4V 7B/13B/34B 03/2024
AVicuna [tang2024avicuna] 100 CLIP ViT MLP, LoRA 1 A6000 GPU Vicuna-v1.5 7B 03/2024
CAT [Ye2024CAT] - ImageBind Linear Projector 1 A100 GPU LLaMA-2 7B 03/2024
Uni-AD [Wang2024Uni-AD] - CLIP ViT-B/32 Video mapping network 8 A100 GPUs LLaMA-2 7B 03/2024
Elysium [wang2024elysium] Varying CLIP-ViT-L T-Selector 24 A100-80G Vicuna - 03/2024
@@
VideoTree [wang2024videotree] Varying EVA-CLIP ViT-G/14 - 4 A6000 GPUs GPT-4 - 04/2024
VTG-LLM [guo2024vtgllm] 96 EVA-CLIP ViT-G/14 Projector 6 ATN 910B LLaMA-2 7B 04/2024
AutoAD III [han2024autoadiii] 8 EVA-CLIP Linear Projector 1 A40 GPU GPT-3.5-turbo - 04/2024
LLaVA-Hound-DPO [zhang2024llavahounddpo] 10 LanguageBind ViT/14 - - GPT-4V - 04/2024
RedViLLM [huang2024redvillm] Varying CLIP ViT-G/14 Temporal Module - Qwen-VL 7B 04/2024
LAVAD [zanella2024lavad] 16 BLIP-2 ViT-L/14 - - LLaMA-2-chat 13B 04/2024
@@
VLM4HOI [bansal2024hoiref] - EVA CLIP Projection Layer 8 V100 GPUs LLaMA-2-Chat 13B 04/2024
@@
Koala [tan2024koala] 64 EVA-CLIP ViT-G/14 Video QFormer+Linear Layer 4 RTX A6000 GPUs Vicuna-v0 7B/13B 04/2024
LongVLM [weng2024longvlm] 100 CLIP ViT-L/14 Projection Layer 4 A100 GPUs Vicuna-v1.1 7B 04/2024
MA-LMM [he2024malmm] 100 EVA-CLIP ViT-G/14 Q-Former 4 A100 GPUs Vicuna 7B 04/2024
@@
MiniGPT4-Video [ataallah2024minigpt4video] Varying EVA-CLIP Projector + LoRA for LLM - LlaMA-2/Mistral 7B 04/2024
MoReVQA [min2024morevqa] 1 FPS - - - PaLM-2 - 04/2024
Pegasus-v1 [jung2024pegasusv1] - Marengo 2.6 VL Alignment Module - GPT-4 - 04/2024
PLLaVA [xu2024pllava] 16 CLIP ViT-L/14 MM Projector - LLaVA 7B/13B/34B 04/2024
ST-LLM [liu2024stllm] Varying CLIP ViT-L/14 Projection Layer 8 A100 GPUs Vicuna-v1.1 7B 04/2024
Tarsier [wang2024tarsier] Varying CLIP ViT MLP 48 A100 GPUs Vicuna 7B/13B 04/2024
TraveLER [shang2024traveler] Varying - - 8 RTX A6000 GPUs GPT-3.5 - 04/2024
V2Xum-LLaMA [hua2024v2xum] 1 FPS CLIP ViT-L/14 Vision Adapter 8 A100 GPUs LLaMA 7B/13B 04/2024
VITRON [fei2024vitron] - - Projection Layer 10 A100 GPUs Vicuna-v1.5 7B 04/2024
X-VARS [held2024xvars] 16 CLIP ViT-L/14 Linear Layer 1 V100 GPU VideoChatGPT - 04/2024
VideoChat2 [li2024mvbench] Varying UMT-L QFormer - Vicuna 7B 05/2024
@@
Shotluck Holmes [luo2024shotluckholmes] 120 SigLip MLP Projection Layer 8 H100 GPUs Vicuna 7B 05/2024
VideoStreaming [qian2024videostreaming] 16 CLIP ViT-L/14 Projecter 32 A100 GPUs Vicuna 7B 05/2024
VideoNarrator [yang2024videonarrator] - CLIP ViT-L/14 Video Projecter 4 A6000 GPUs Baichuan 7B 05/2024
TOPA [li2024topa] 10 CLIP ViT-L/14 Linear Projector 4 A100 GPUs Llama2 7/8/13B 05/2024
MotionLLM [chen2024motionllm] 8 LanguageBind/VQ-VAE Linear Layer A100 GPUs Vicuna 7B 05/2024
Artemis [qiu2024artemis] Varying CLIP ViT-L/14 MLP 8 A800 GPUs Vicuna 7B 06/2024
DrVideo [ma2024drvideo] 2 FPS CLIP ViT-L/14 - - LaViLa 7B 06/2024
@@
EmoLLM [yang2024emollm] 8 CLIP ViT-L/14 Multi-Perspective Visual Projection 4 4090 GPUs Vicuna-v1.5 7B 06/2024
@@
FTFV-LLM [chen2024ftfvllm] 8 OpenCLIP ViT-G/14 Vision-Language Adapter 64 A100 GPUs Vicuna 7B 06/2024
Flash-VStream [zhang2024flashvstream] 8 - MLP 8 A100 GPUs GPT-3.5 - 06/2024
@@
Holmes-VAD [zhang2024holmesvad] Varying CLIP ViT-L/14 Temporal Sampler 2 A100 GPUs Vicuna-v1.5 7B 06/2024
LongVA [zhang2024longva] 1 FPS CLIP ViT-L/14 MLP 8 A100 GPUs Mistral-7B-Instruct-v0.2 7B 06/2024
OmAgent [zhang2024omagent] Varying - - - GPT-4o - 06/2024
GIT-LLaVA [kalarani2024gitllava] 6 CLIP ViT-L/14 MLP 4 A100 GPUs Vicuna 7B 06/2024
@@
ShareGPT4Video [chen2024sharegpt4video] 16 CLIP ViT-L/14 Linear Layer + LoRA for LLM 8 A100 GPUs LLaMA-3 8B 06/2024
LVNet [park2024lvnet] - - - - GPT-4 - 06/2024
VIM [du2024vim] - EVA-CLIP ViT-G/14 Q-former+LoRA in LLM 8 A100 GPUs Vicuna-v1.1 13B 06/2024
Video-SALMONN [sun2024videosalmonn] 2 FPS BLIP-2 ViT-G/14 MRC Q-Former - GPT-4 - 06/2024
VideoGPT+ [maaz2024videogpt+] 16/8 CLIP ViT-L/14 Linear Layer 8 A100 GPUs Phi-3-Mini 3.8B 06/2024
Vriptor [yang2024vript] - EVA-CLIP ViT-G/14 - A100 GPUs - - 06/2024
MM-Screenplayer [wu2024mmscreenplayer] - - - - GPT-4-turbo - 06/2024
VideoLLaMA 2 [cheng2024videollama2] 16 CLIP ViT-L/14 Linear Layer +STC connector - LLaVA-1.5 7B 06/2024
@@
VideoLLM-online [chen2024videollmonline] 2 FPS CLIP ViT-L/14 MLP projector + LoRA for LLM - Llama-2-Chat/Llama-3-Instruct 7B/8B 06/2024
@@ III-B2 Fine-tuning Vid-LLMs
  • LLM Fully Fine-tuning: This fine-tuning method does not use any adapters but instead employs supervised training with a lower learning rate, updating all the parameters in the LLM. This method allows the Vid-LLM to fully adapt to the respective task and achieve good performance, especially when the target task is quite different from the pretraining tasks. For end-to-end Vid-LLMs, especially those in the Video Embedder $\times$ LLM category, the Video Embedder may also be fine-tuned for more comprehensive learning. However, this method consumes more computational resources than adapter-based fine-tuning methods and may potentially impair the inherent capabilities of the LLM, such as zero-shot and in-context learning. Vid-LLM adopted LLM Fully Fine-tuning include AV-LLM [Shu2023AV-LLM] and Vid2Seq [yang2023vid2seq]. In [Shu2023AV-LLM], there are both fully fine-tuning and adapter fine-tuning versions of Vid-LLMs, and the former’s performance is better than the latter.
  • Connective Adapter Fine-tuning: Here, the term “Connective” refers to adapters that bridge the Video Embedder and the LLM externally, enabling information from the video to flow into the LLM through the Connective Adapter. As illustrated in Figure 6, during training, the parameters of both the Video Embedder and the LLM are frozen, and only the parameters of the Connective Adapter are updated. Common Connective Adapters include MLP/Linear Layer and Q-former [li2023blip], their combinations, etc., whose primary function is to map video embeddings from the visual semantic space to the text semantic space of the LLM input tokens (i.e., modality alignment). Typically, fine-tuning only the Connective Adapter does not alter the LLM’s inherent behavior.
@@ IV-A Closed-ended Evaluation
Closed-ended evaluations use pre-defined answers or structured formats [yin2023survey]. These include multiple-choice questions (TVQA [lei2018tvqa], How2QA [How2QA], STAR [wu2021star]) and questions with structured formats for direct comparison with ground truth (MSRVTT-QA [xu2017video], MSVD-QA [xu2017video], MVBench [li2024mvbench]). Multiple-choice performance is evaluated via accuracy percentages, while structured formats use metrics like CIDEr [vedantam2015cider], METEOR [banerjee2005meteor], ROUGE [lin2004rouge], and SPICE [anderson2016spice] to compare predictions with ground truth. Notable benchmarks include MSRVTT-QA [xu2017video], TVQA [lei2018tvqa], MVBench [li2024mvbench], EgoSchema [mangalam2023egoschema], and Video-MME [fu2024video-mme]. Each targets different video understanding aspects: TGIF-QA [jang2017tgif] focuses on action recognition, state transition, frame-level QA, and counting; ActivityNet-QA [yu2019activitynet] covers motion, spatial, temporal, and free-form dimensions; VidComposition [tang2024vidcomposition] emphasizes compositional reasoning; while NExT-QA [NExT-QA] and MLVU [zhou2024mlvu] include causal and temporal action reasoning. These diverse question types test various reasoning abilities, though many benchmarks still exhibit domain biases toward common scenarios and lack diversity in rare events or unusual contexts.
@@
MSRVTT-QA [xu2017video] 2,990 2,990 15.2 Open-domain 72,821 Closed-ended & open-ended (what/who/how/when/where)
MSVD-QA [xu2017video] 504 504 9.8 Open-domain 50,505 Closed-ended & open-ended (what/who/how/when/where)
TGIF-QA [jang2017tgif] 9,575 9,575 3.0 Open-domain 8,506 Closed-ended & open-ended, action, transition, counting
ActivityNet-QA [yu2019activitynet] 800 800 111.4 Human activity 8,000 Closed-ended & open-ended (what/who/how/when/where/why)
TVQA [lei2018tvqa] 2,179 15,253 11.2 TV show 15,253 Closed-ended, multiple-choice
How2QA [How2QA] 1,166 2,852 15.3 TV episode 2,852 Multiple-choice
STAR [wu2021star] 914 7,098 11.9 Human action 7,098 Multiple-choice
NExT-QA [NExT-QA] 1,000 1,000 39.5 Daily life 8,564 Multiple-choice and open-ended (causal, temporal, descriptive)
MVBench [li2024mvbench] 3,641 3,641 16.0 Open-domain 4,000 Closed-ended (various tasks)
Video-Bench [ning2023video] 5,917 5,917 56.0 Open-domain 17,036 Open-ended (various tasks)
EgoSchema [mangalam2023egoschema] 5,063 5,063 180.0 Egocentric activity 5,063 Closed-ended, procedural understanding
AutoEval-Video [chen2023autoeval] 327 327 14.6 Open-domain 327 Open-ended evaluation
TempCompass [liu2024tempcompass] 410 500 11.4 Open-domain 7,540 Both closed-ended & open-ended, temporal reasoning
Video-MME [fu2024video-mme] 900 900 1,017.9 Open-domain 2,700 Closed-ended evaluation
VideoVista [li2024videovista] 894 3,400 131.0 Open-domain 25,000 Open-ended (descriptive, causal, predictive)
CinePile [rawal2024cinepile] 9,396 9,396 160.0 Movie 303,828 Movie understanding, open-ended
SOK-Bench [wang2024sok] 10,000 10,000 - Open-domain 44,000 Open-ended, subject-oriented knowledge
SFD [ghermi2024short] 1,078 1,078 780.0 Movies 4,885 Multiple-choice, open-ended
EditVid-QA [xu2024beyond] 32,000 32,000 - Entertainment 204,000 Open-ended, editing techniques
InfiniBench [ataallah2024infinibench] 1,219 1,219 4,460.4 Movie/TV show 108,200 Closed-ended & open-ended, long-form understanding
MLVU [zhou2024mlvu] 1,334 1,334 180.0-7,200.0 Open-domain 2,593 Both closed-ended and open-ended, long-form understanding
MMWorld [he2024mmworld] 1,910 1,910 102.3 Open-domain 1,599 Open-ended, multimodal evaluation
VELOCITI [saravanan2024velociti] 900 900 10.0 Movie - Open-ended, visual understanding
VidComposition [tang2024vidcomposition] 982 982 - Movie/TV show 1,706 Multiple-choice, compositional reasoning
EAGLE [bi2024eagle] 36,000 36,000 16.0-76.0 Egocentric activity 400,000 Open-ended, procedural understanding